The following table shows the relationship between cigarette smoking and lung cancer among male cases and controls.

Cases

Controls

Cigarette smoker

1,350

1,296

Non-Smoker

7

61

Total

1,357

1,357

Table 1. Smoking status before onset of the present illness, lung cancer cases and matched controls with other diseases, Great Britain, 1948-1952.

Question 11 (2 points) Calculate the odds of smoking among the cases?
Answer #1

According to the given data

Total Cigarette smoker cases are 1350

Total non smoker cases are 7

Total cases are 1357

Among the cases total odd cases =?

Odds ratio = probability/ (1-probability) = proportion /(1-proportion)

Therefore (1,350/1,357) / (7/1,357) = 1,350/7 = 192.9:1

Patients in case group is 193 times that would probably smoke